OpenStudy (anonymous):

Evaluate the combination. 5C4 A.) 1 B.) 4 C.) 5 D.) 10

OpenStudy (anonymous):

the answer is a.1

OpenStudy (anonymous):

Thank youu!

OpenStudy (anonymous):

ur welcome sweethart

OpenStudy (nurali):

\[_{n}C _{r}=\frac{ n! }{ r!(n-r)! }\] Where n=5 and r=4 \[5C4=\frac{ 5! }{ 4!(5-4)!}\] \[5C4=\frac{ 5! }{ 4!(1)! }\] \[5C4=\frac{ 5*4*3*2*1 }{ 4*3*2*1*(1) }\]  \[5C4=5\]

OpenStudy (nurali):

@snowman97 answer is wrong.